# The effect of self-induced Marangoni flow on polar-nematic waves in active-matter systems

### v1

@@ 3 Marangoni flow-induced change in polar-nematic waves
[…]
Figure 1: Snapshots of fully developed polar-nematic travelling waves in 2D for different values of the Marangoni parameter illustrate the effect of Marangoni flow. Shown are (a) $\text{Ma}=0$, (b) $\text{Ma}=10$, (c) $\text{Ma}=30$, and (d) $\text{Ma}=150$. Accompanying movies in the Supplementary Material further illustrate the wave dynamics. The mean density is $c_{0}=0.09$, the diffusivity is $d=0.1$, and a $800\times 800$ domain with periodic boundaries is used. The background coloring indicates the number density $c$ as indicated in the color bar. Black arrows indicate the vectorial polar order parameter field $(\text{Re}(\rho_{1}),\text{Im}(\rho_{1}))$. No stable waves are found for $\text{Ma}>160$. All remaining parameters are given in the main text.
@@ 4.1 Uniform states and their linear stability
First, we discuss the uniform steady states of Eqs. (4) that exist at the present parameters [Eqs. (2)] using the mean density $c_{0}$ as control parameter. The trivial state $\rho_{1}^{\mathrm{t}}=\rho_{2}^{\mathrm{t}}=0$ exists for all $c_{0}$. Two other branches of uniform steady states exist in certain ranges of $c_{0}$: The first branch bifurcates in a pitchfork bifurcation from the trivial state and exists for all $c_{0}>\beta_{0}/\beta_{1}=0.074$. It corresponds to purely nematic uniform states
@@
$\rho_{1}^{\mathrm{pn}}$ $=$ ±$\pm\sqrt{(\beta_{2}-\beta_{4}\rho_{2}^{\mathrm{pn}})^{-1}[\beta_{3}(\rho_{2}^{\mathrm{pn}})^{3}+(\beta_{0}-c_{0}\beta_{1})\rho_{2}^{\mathrm{pn}}]},$ (10)
Fig. 2 presents a bifurcation diagram featuring all uniform steady states. Both, polar order parameter $\rho_{1}$ and nematic order parameter $\rho_{2}$, are given as a function of $c_{0}$. The three-dimensional representation together with the projections of the bifurcation diagram onto the ($c_{0},\rho_{1}$)- and ($c_{0},\rho_{2}$)-planes well represent the connections between the different states and the symmetries of the system.
@@
Figure 3: (a) Linear stability of the purely nematic uniform state $\rho_{1,2}^{\mathrm{n}}$ [Eq. (9)] in the plane spanned by mean density $c_{0}$ and Marangoni parameter Ma. Labels “s”, “w” and “m” mark the linearly stable region, the region of the wave instability and the region of mixed wave-uniform instability, respectively, as described in the main text. Panels (b) and (c) give typical dispersion relations, i.e. growth rates $\text{Re}(\lambda)$ as a function of the wavenumber $k$, in region w at $\text{Ma}=0$, $c_{0}=0.09$ and in region m at $\text{Ma}=0$, $c_{0}=0.095$, respectively. Dashed (solid) lines correspond to complex (real) eigenvalues $\lambda$. (d) Linear stability of the mixed polar-nematic uniform state $\rho_{1,2}^{\mathrm{pn}}$ [Eq. (4.1)]. In the regions marked by I [II] at most one [two] eigenvalues have a positive real part for any given $k$. Panels (e) and (f) give dispersion relations in region I at $\text{Ma}=0$, $c_{0}=0.12$ and region II at $\text{Ma}=0$, $c_{0}=0.15$, respectively.
@@
Note that the Jacobian matrix $\mathbf{\underline{J}}$ is complex due to the term $ik\mathbf{\underline{J}}_{1}$ that is related to self-propulsion. In consequence, for any wave number $k$, up to three different eigenvalues may exist, each either complex or real. More specifically, in the case of the trivial uniform state $\rho_{1,2}^{\mathrm{s}}=\rho_{1,2}^{\mathrm{t}}=0$ and in the case of the purely nematic state $\rho_{1,2}^{\mathrm{s}}=\rho_{1,2}^{\mathrm{n}}$ [Eq. (9)] the resulting characteristic polynomial $\det(\mathbf{\underline{J}}-\lambda\mathbb{I})$ has only real coefficients. Therefore, all three eigenvalues are either real, or appear in complex conjugate pairs. This feature is similar to isotropic systems where the Jacobian is always real. In contrast, for the mixed polar-nematic state with $\rho_{1,2}^{\mathrm{s}}=\rho_{1,2}^{\mathrm{pn}}$ the characteristic polynomial has complex coefficients resulting in three different, in general, complex eigenvalues.
@@
The eigenvalue problem Eq. (12) is solved numerically for various mean densities $c_{0}$ and Marangoni parameters Ma, to determine the regions of linear stability (where $\text{Re}(\lambda)<0$) for the three uniform steady states. Resulting stability diagrams for the nematic and the nematic-polar states in the plane $(c_{0},\text{Ma})$ are given in Figs. 3 (a) and 3 (d), respectively, as explained in the following. As expected, the trivial state $\rho_{1}^{\mathrm{t}}=\rho_{2}^{\mathrm{t}}=0$ is (independently of Ma) linearly stable at small $c_{0}$ and becomes unstable at $c_{0}=\beta_{0}/\beta_{1}=0.074$. The instability is stationary and large-scale, i.e. the onset is at $k=0$. At the instability threshold, the above-mentioned pitchfork bifurcation occurs where the uniform purely nematic state $\rho_{1,2}^{\mathrm{n}}$ [Eq. (9)] emerges. For slightly supercritical $c_{0}$, the leading eigenvalue of the trivial state is real and is given by $\lambda(k)=-\beta_{0}+c_{0}\beta_{1}-(d+\gamma_{2})k^{2}-k^{2}/[4(\alpha_{0}+c_{0}(\alpha_{1}+\beta_{1})-\beta_{0})]+O(k^{4})$.
@@
Following the $\rho_{1,2}^{\mathrm{n}}$-branch for $\text{Ma}=0$ when gradually increasing $c_{0}$, we find a wave instability at $c_{0}=0.086$, i.e. a small-scale oscillatory instability, sometimes also called a finite-wavelength Hopf or oscillatrory Turing instability. It is marked by a blue square symbol and the letter “W” in Fig. 2. Beyond this point the uniform nematic state is always unstable for sufficiently large domain sizes. At the wave bifurcation travelling nematic bands emerge as discussed in Ref. [14]. Fig. 3 (b) shows a typical dispersion relation in the range between the wave instability and the pitchfork bifurcation where the uniform $\rho_{1,2}^{\mathrm{pn}}$-state emerges from $\rho_{1,2}^{\mathrm{n}}$. Dashed (solid) lines correspond to complex (real) eigenvalues $\lambda$. The dispersion relations are qualitatively similar in the entire region of Fig. 3 (a) that is marked “w”.
@@
At the onset of the wave instability (W-point in Fig. 2), the leading eigenvalue of the Jacobian is purely imaginary at some critical wave number $k_{c}$, i.e. $\lambda(k_{c})=i\Omega$, where $V=\Omega/k_{c}$ is the speed of the propagating non-decaying wave that is at onset of infinitesimal amplitude. Further increasing $c_{0}$ beyond the W-point, the already unstable uniform nematic state at $c_{0}=0.091$ additionally undergoes the above described spatially uniform stationary instability where the $\rho_{1,2}^{\mathrm{pn}}$-branch emerges. Beyond this point, the dispersion relation for the $\rho_{1,2}^{\mathrm{n}}$-state features a real positive eigenvalue at small $k$ that connects to a pair of complex-conjugate eigenvalues at larger $k$, see Fig. 3 (c). Together they form a continuous band of unstable wavenumbers. Note that in the vicinity of $c_{0}=0.091$ the wavenumber band of the stationary instability becomes very narrow. Exactly at $c_{0}=0.091$ it only contains the point $k=0$ implying that a this single parameter value the dispersion relation corresponds to a conserved-Hopf instability [19, 40].
@@
We call the combination of real and complex eigenvalues above $c_{0}=0.091$“mixed mode” and mark the corresponding region in Fig. 3 (a) by “m”. Up to here, our discussion of the wave instability has focused on the dry case with Ma$=0$. Inspecting Fig. 3 (a) we see that an inclusion of Marangoni flow suppresses the wave instability of the nematic state. It completely disappears at $\text{Ma}\approx 100$. Interestingly, this further implies that in contrast to the case studied in [40], here the conserved-Hopf instability appears as a codimension-two instability exactly where the line separating the s- and w-region in Fig. 3 (a) ends.
@@
It was shown already in Ref. [14] for the dry case that the time evolution of the system in the mixed mode region is dominated by nematic-polar wave patterns characterised by spatially distributed polar and nematic order parameter fields. Analysing the linear stability of the uniform polar-nematic states $\rho_{1,2}^{\mathrm{pn}}$ [Eq. (4.1)] for a spatially extended system we find as expected that the state inherits the wave instability from the $\rho_{1,2}^{\mathrm{n}}$-branch it emerges from at $c_{0}=0.091$. A typical dispersion relation at moderate $c_{0}$ is given in Fig. 3 (e), and shows that the band of unstable wavenumbers is bound away from zero. However, at lower and higher densities the unstable band can also start at $k=0$ (not shown). The $\rho_{1,2}^{\mathrm{pn}}$-branch retains this instability till it ends at $c_{0}=0.2425$. It even acquires a second instability mode in an intermediate $c_{0}$-range, see the dispersion relation in Fig. 3 (f). We mark the regions where one and two wave modes are unstable in Fig. 3 (d) by “I” and “II”, respectively.
@@
Moving again from the dry to the wet case, we see in the stability diagram Fig. 3 (d) that an increase in the Marangoni parameter tends to suppress the wave instabilities: At $\text{Ma}\approx 100$ the II-region has been replaced by I-regions, i.e. the second wave mode has been suppressed, and a stable range of concentrations $c_{0}$ marked by “s” in Fig. 3 (d) has started to develop around $c_{0}=0.2$. Upon inspecting the eigenvectors that correspond to the maximum of the dispersion curves for two unstable modes in Fig. 3 (f), we find that the slow growing mode is a slightly asymmetric polar-nematic wave with the amplitude of the field $\rho_{2}$ being smaller than the amplitudes of fields $c$ and $\rho_{1}$. On the contrary, the fastest growing mode corresponds to a fully symmetric polar-nematic wave with the amplitudes of all three fields $c$ and $\rho_{1,2}$ being almost identical.
@@
Note the codimension-2 point that occurs where in Fig. 3 (d) the two lines meet that separate regions I and II. There, the range of linearly stable concentrations starts, i.e. with other words, the two wave instabilities have a simultaneous onset. Adjusting parameters that control the corresponding critical wavenumbers one might in the future be able to find resonances with interesting nonlinear behaviour.
@@ 4.2 Travelling waves in the dry system
The dispersion relations in Figs. 3 (e) and 3 (f) indicate that in regions I and II of Fig. 3 (d), respectively, one and two wave modes are unstable. As both modes correspond to bands of unstable wavenumbers bound away from zero, the one-mode [two-mode] case corresponds to two [four] bifurcations where branches of travelling waves emerge when using the domain size as control parameter. In situations as in the first row of Fig. 3 there is one wave mode with two [Fig. 3 (b)] or one [Fig. 3 (c)] bifurcations. Depending on the specific parameter values the bifurcations may be super- or subcritical, the branches may be connected or independent – this is determined by the fully nonlinear behaviour. Directly at each bifurcation there exist harmonic waves of infinitesimal amplitude with propagation speed $V_{c}=\text{Im}(\lambda)/k_{c}$ and wavelength $L_{c}=2\pi/k_{c}$, where $k_{c}$ corresponds to the neutral mode, i.e. $\text{Re}(\lambda)(k_{c})=0$. To gain knowledge about the branches of fully nonlinear travelling waves one has to resort to numerical methods.
[…]
@@
First, we consider the dry system ($\text{Ma}=0$) at fixed mean density $c_{0}=0.09$ (where the polar-nematic state does not yet exist) and examine the wave bifurcation where a branch of travelling waves emerges from the uniform nematic state. As determined in section 4.1, for $c_{0}>0.086$ this uniform state is unstable in a certain range of wavenumbers that depends on $c_{0}$. When we fix $c_{0}$ and use the domain size $L$ as control parameter, i.e. as principal continuation parameter, we encounter two disconnected branches of travelling waves that emerge where the dispersion relation in Fig. 3 (b) has its two zero crossings. Fig. 4 shows in panel (a) the propagation speed $V$, in panel (b) the amplitude of the density field $A=\text{max}(c)-\text{min}(c)$, and in panel (c) the polar and mean nematic order parameter fields $\langle\rho_{1,2}\rangle=L^{-1}\int_{0}^{L}\rho_{1,2}(\chi),\,d\chi$ as a function of the domain size $L$. The wave with the larger [smaller] amplitude corresponds to a travelling density elevation [depression] both propagating in the uniform background, as shown in the insets of Fig. 4. When numerically integrating Eqs. (4) in time using a naive continuation scheme as described in section 3, at all $L$ the system converges to the state with the elevation (density hump, black line), as indicated by small circles in Fig. 4 (a) and (b). The travelling holes (red line) are never found in time simulations, which indicates that they are either unstable or linearly stable but with a very small basin of attraction.
@@
Second, we look at the dry system at the larger mean density $c_{0}=0.12$ where also the uniform polar-nematic state exists. The linear stability results then indicate that overall three types of small-amplitude travelling wave states may emerge from the uniform states (one from purely nematic and two from mixed polar-nematic). The corresponding zero crossings of the dispersion relations (Fig. 3) correspond to three wave bifurcations and may give rise to up to three disconnected branches of finite-amplitude waves (the exact number depends on the fully nonlinear behaviour). At $c_{0}=0.12$ we find three disconnected branches, see in Figs. 5 (a)-(c). The first one bifurcates at the critical $k$ from the purely nematic state (red dashed line) and the other two at the critical $k$’s for the mixed polar-nematic state (solid black and dot-dashed blue line). The dot-dashed branch undergoes two saddle-node bifurcations, however, at large domain sizes three clearly distinguished travelling waves remain. They correspond to a single-hump wave (solid black line) as shown in Fig. 5(1,3), a double-hump wave (blue dot-dashed line) as shown in Fig. 5(2), and a travelling-hole state (red dashed line) as shown in Fig. 5(4). The double-hump state corresponds to a large travelling density cluster followed at a small distance by a small satellite peak. Direct numerical time integration of Eqs. (4) using the naive continuation algorithm as described above only converges to single-hump waves (as indicated by star symbols in 5 (a)-(c)) indicating their linear stability. The other types of travelling waves were not found in the time simulations.
Figure 5: Three disconnected branches of finite-amplitude travelling waves for a dry system (Ma$=0$) at $c_{0}=0.12$ and $d=0.01$. (a) Wave speed $V$, (b) amplitude of the density field $A=\text{max}(c)-\text{min}(c)$ and (c) average nematic order parameter field $\langle\rho_{2}\rangle$ as a function of domain size $L$. The four red filled circles marked (1) to (4) in panel (b) indicate loci of density profiles shown in the bottom panels. Star symbols indicate stable wave states obtained by direct numerical time integration of Eqs. (4).

### v2

@@ 3 Marangoni flow-induced change in polar-nematic waves
As shown in Ref. [14], the dry system, i.e. without the presence of Marangoni flow, develops polar-nematic travelling waves for mean densities above $c_{0}=\beta_{0}/\beta_{1}=0.074$. Before embarking on a detailed analysis, we first demonstrate that self-induced Marangoni flow influences the propagation of these density waves. We use the long-wave variant [Eq. (3)] to calculate the Marangoni flow [33] and numerically solve Eqs. (6) in 2D using central differences for spatial derivatives and a semi-implicit second-order Heun method for time integration. As initial condition, we use a uniform disordered state with small-amplitude random initial perturbations with a slight polar and nematic bias in the $x$-direction. Eqs. (6) are numerically integrated over a sufficiently long time interval of $5000$ units at $\text{Ma}=0$ to allow the system to reach a steady state. Then the average wave number associated with the Fourier-transformed density field $\int|\mathbf{k}|\hat{c}(\mathbf{k},t)\,d\mathbf{k}$ fluctuates about a constant value and does not show any trend as a function of time. The density field $c$ recorded over the subsequent time interval $\Delta t=1000$ is used to extract the statistics of the speed of the waves (not shown). The dynamics of the system at $\text{Ma}>0$ is obtained using naive parameter continuation. Namely, after the system is integrated for $\Delta t=6000$, the Marangoni parameter Ma is abruptly incremented by $10$ and Eqs. (6) are further integrated without setting a new initial condition.
Figure 2: Snapshots of fully developed polar-nematic travelling waves in 2D for different values of the Marangoni parameter illustrate the effect of Marangoni flow. Shown are (a) $\text{Ma}=0$, (b) $\text{Ma}=10$, (c) $\text{Ma}=30$, and (d) $\text{Ma}=150$. Accompanying movies in the Supplementary Material further illustrate the wave dynamics. The mean density is $c_{0}=0.09$, the diffusivity is $d=0.1$, and a $800\times 800$ domain with periodic boundaries is used. The background colouring indicates the number density $c$ as indicated in the color bar. Black arrows indicate the vectorial polar order parameter field $(\text{Re}(\rho_{1}),\text{Im}(\rho_{1}))$. No stable waves are found for $\text{Ma}>160$. All remaining parameters are given in the main text.
@@ 4.1 Uniform states and their linear stability
First, we discussing the uniform steady states of Eqs. (4) that exist at the present parameters [Eqs. (2)] using the mean density $c_{0}$ as control parameter. The trivial state $\rho_{1}^{\mathrm{t}}=\rho_{2}^{\mathrm{t}}=0$ exists for all $c_{0}$. Two other branches of uniform steady states exist in certain ranges of $c_{0}$: The first branch bifurcates in a pitchfork bifurcation from the trivial state and exists for all $c_{0}>\beta_{0}/\beta_{1}=0.074$. It corresponds to purely nematic uniform states
@@
$\rho_{1}^{\mathrm{pn}}$ $=$ $\pm\sqrt{(\beta_{2}-\beta_{4}\rho_{2}^{\mathrm{pn}})^{-1}[\beta_{3}(\rho_{2}^{\mathrm{pn}})^{3}+(\beta_{0}-c_{0}\beta_{1})\rho_{2}^{\mathrm{pn}}]},$ (10)
Note that Eqs. (9),4.1) are valid for dry systems as well as for wet systems, since the Marangoni flow induced by a uniform density distribution vanishes. Fig. 3 presents a bifurcation diagram featuring all uniform steady states. Both, polar order parameter $\rho_{1}$ and nematic order parameter $\rho_{2}$, are given as a function of $c_{0}$. The three-dimensional representation together with the projections of the bifurcation diagram onto the ($c_{0},\rho_{1}$)- and ($c_{0},\rho_{2}$)-planes well represent the connections between the different states and the symmetries of the system. In the case of a dry system, the bifurcation diagram in Fig. 3 can be directly compared with a slice of the stability diagram obtained in Ref. [14] for varying $c_{0}$ and $\alpha_{2}$. Indeed, at the value $\alpha_{2}=1.5$ here used the critical density for the nematic-polar transition is $c_{0}=0.091$ as obtained in Ref. [14].
@@
Figure 4: (a) Linear stability of the purely nematic uniform state $\rho_{1,2}^{\mathrm{n}}$ [Eq. (9)] in the plane spanned by mean density $c_{0}$ and Marangoni parameter Ma. Labels “s”, “w” and “m” mark the linearly stable region, i.e., the region where all wave numbers are stable, the region of the wave instability (oscillatory instability of a finite band of wave numbers bound away from zero) and the region of mixed mode instability (unstable band of wave numbers $0\leq k<k_{c}$ with real or complex eigenvalues depending on $k$), respectively, see main text. Panels (b) and (c) give typical dispersion relations, i.e., growth rates $\text{Re}(\lambda)$ for the leading eigenvalues as a function of the wavenumber $k$, in region w at $c_{0}=0.09$ and in region m at $c_{0}=0.095$, respectively, at two different values of Ma as indicated near each curve. Dashed (solid) lines correspond to complex (real) eigenvalues $\lambda$. (d) Linear stability of the mixed polar-nematic uniform state $\rho_{1,2}^{\mathrm{pn}}$ [Eq. (4.1)]. In the regions marked by I [II] at most one [two] eigenvalues have a positive real part for any given $k$. Panels (e) and (f) give dispersion relations in region I at $\text{Ma}=0$, $c_{0}=0.12$ and region II at $\text{Ma}=30$, $c_{0}=0.18$, respectively.
@@
Note that the Jacobian matrix $\mathbf{\underline{J}}$ is complex due to the term $ik\mathbf{\underline{J}}_{1}$ that is related to self-propulsion. In consequence, for any wave number $k$, up to three different eigenvalues may exist, each either complex or real. More specifically, in the case of the trivial uniform state $\rho_{1,2}^{\mathrm{s}}=\rho_{1,2}^{\mathrm{t}}=0$ and in the case of the purely nematic state $\rho_{1,2}^{\mathrm{s}}=\rho_{1,2}^{\mathrm{n}}$ [Eq. (9)] the resulting characteristic polynomial $\det(\mathbf{\underline{J}}-\lambda\mathbb{I})$ has only real coefficients. Therefore, all three eigenvalues are either real, or one eigenvalue is real and the remaining two form a complex conjugate pair. This feature is similar to isotropic systems where the Jacobian is always real. In contrast, for the mixed polar-nematic state with $\rho_{1,2}^{\mathrm{s}}=\rho_{1,2}^{\mathrm{pn}}$ the characteristic polynomial has complex coefficients resulting in three different, in general, complex eigenvalues.
@@
The eigenvalue problem Eq. (12) is solved numerically for various mean densities $c_{0}$ and Marangoni parameters Ma, to determine the regions of linear stability (where $\text{Re}(\lambda)<0$) for the three uniform steady states. Resulting stability diagrams for the nematic and the nematic-polar states in the plane $(c_{0},\text{Ma})$ are given in Figs. 4 (a) and 4 (d), respectively, as explained in the following. As expected, the trivial state $\rho_{1}^{\mathrm{t}}=\rho_{2}^{\mathrm{t}}=0$ is (independently of Ma) linearly stable at small $c_{0}$ and becomes unstable at $c_{0}=\beta_{0}/\beta_{1}=0.074$. The instability is stationary and large-scale, i.e. the onset is at $k=0$. At the instability threshold, the above-mentioned pitchfork bifurcation occurs where the uniform purely nematic state $\rho_{1,2}^{\mathrm{n}}$ [Eq. (9)] emerges. For slightly supercritical $c_{0}$, the leading eigenvalue of the trivial state is real and is given by $\lambda(k)=-\beta_{0}+c_{0}\beta_{1}-(d+\gamma_{2})k^{2}-k^{2}/[4(\alpha_{0}+c_{0}(\alpha_{1}+\beta_{1})-\beta_{0})]+O(k^{4})$.
@@
Following the $\rho_{1,2}^{\mathrm{n}}$-branch for $\text{Ma}=0$ when gradually increasing $c_{0}$, we find a wave instability at $c_{0}=0.086$, i.e. a small-scale oscillatory instability, sometimes also called a finite-wavelength Hopf or oscillatory Turing instability. It is marked by a blue square symbol and the letter “w” in Fig. 3. Beyond this point the uniform nematic state is always unstable for sufficiently large domain sizes. At the wave bifurcation travelling nematic bands emerge as discussed in Ref. [14]. Fig. 4 (b) shows the dispersion relation, represented by the real parts of the two leading eigenvalues in the parameter range between the onset of the wave instability and the pitchfork bifurcation where the uniform $\rho_{1,2}^{\mathrm{pn}}$-state emerges from $\rho_{1,2}^{\mathrm{n}}$. Dashed (solid) lines correspond to complex (real) eigenvalues $\lambda$. The dispersion relations are qualitatively similar in the entire region marked “w” in Fig. 4 (a). In this region, there exists a band of unstable wave numbers $k$, namely, the eigenvalue $\lambda$ is complex with a positive real part. The Marangoni flow suppresses the instability as shown in Fig. 4 (b).
@@
At the onset of the wave instability (W-point in Fig. 3), the leading eigenvalue of the Jacobian is purely imaginary at some critical wave number $k_{c}$, i.e. $\lambda(k_{c})=i\Omega$, where $V=\Omega/k_{c}$ is the speed of the propagating non-decaying wave that is at onset of infinitesimal amplitude. Further increasing $c_{0}$ beyond the W-point, the already unstable uniform nematic state at $c_{0}=0.091$ additionally undergoes the above described spatially uniform stationary instability where the $\rho_{1,2}^{\mathrm{pn}}$-branch emerges. Beyond this point, the dispersion relation for the $\rho_{1,2}^{\mathrm{n}}$-state features a real positive eigenvalue at small $k$ that connects to a pair of complex conjugate eigenvalues at larger $k$, see Fig. 4 (c). Together they form a continuous band of unstable wave numbers. Note that in the vicinity of $c_{0}=0.091$ the wavenumber band of the stationary instability becomes very narrow. Exactly at $c_{0}=0.091$ it only contains the point $k=0$ implying that a this single parameter value the dispersion relation corresponds to a conserved-Hopf instability [19, 43].
@@
We call the combination of real and complex eigenvalues above $c_{0}=0.091$ “mixed mode” and mark the corresponding region in Fig. 4 (a) by “m”. Here, large-scale perturbations are monotonically unstable, while perturbations of smaller scale are oscillatory unstable. However, both mode types form a common band of unstable wave numbers. Sometimes the transition point between real and complex modes within the band is called an “exceptional point.” Note however that such points only directly affect the primary bifurcations if they occur directly at zero crossing - then they correspond to a Takens-Bogdanov bifurcation. As they are detected in a linear analysis their potential influence on the nonlinear behaviour needs additional analysis. Inspecting Fig. 4 (a) we see that an inclusion of Marangoni flow suppresses the wave instability of the nematic state. It completely disappears at $\text{Ma}\approx 100$. Interestingly, this further implies that in contrast to the case studied in [43], here the conserved-Hopf instability appears as a codimension-2 instability exactly where the line separating the s- and w-region in Fig. 4 (a) ends.
@@
It was shown already in Ref. [14] for the dry case that the time evolution of the system in the mixed mode region is dominated by nematic-polar wave patterns characterised by spatially distributed polar and nematic order parameter fields. Analysing the linear stability of the uniform polar-nematic states $\rho_{1,2}^{\mathrm{pn}}$ [Eq. (4.1)] for a spatially extended system we find as expected that the state inherits the wave instability from the $\rho_{1,2}^{\mathrm{n}}$-branch it emerges from at $c_{0}=0.091$. A typical dispersion relation at moderate $c_{0}$ is given in Fig. 4 (e), and shows that the band of unstable wave numbers is bound away from zero. However, at lower and higher densities the unstable band can also start at $k=0$ (not shown).
The $\rho_{1,2}^{\mathrm{pn}}$-branch retains this instability till it ends at $c_{0}=0.2425$. It even acquires a second instability mode in an intermediate $c_{0}$-range, see the dispersion relation in Fig. 4 (f) shown for $c_{0}=0.18$ and $\text{Ma}=30$ close to the codimension-2 point in Fig. 4 (d), where the lines separating the regions “I” and “II” intersect. In these regions I and II there exist one and two wave number bands unstable wave modes, respectively.
@@
From the stability diagram Fig. 4 (d) we see that an increase in the Marangoni parameter tends to suppress the wave instabilities: At $\text{Ma}\approx 100$ the II-region has been replaced by I-regions, i.e. the second wave mode has been suppressed, and a stable range of concentrations $c_{0}$ marked by “s” in Fig. 4 (d) has started to develop around $c_{0}=0.2$.
@@
Note at the codimension-2 point in Fig. 4 (d), where the two lines meet that separate regions I and II, the two wave instabilities have a simultaneous onset. Adjusting parameters that control the corresponding critical wave numbers one might in the future be able to find resonances with interesting nonlinear behaviour.
@@ 4.2 Travelling waves in the dry system
The dispersion relations in Figs. 4 (e) and 4 (f) indicate that in regions I and II of Fig. 4 (d), respectively, one and two wave modes are unstable. As both modes correspond to bands of unstable wave numbers bound away from zero, the one-mode [two-mode] case corresponds to two [four] bifurcations where branches of travelling waves emerge when using the domain size as control parameter. In situations as in the first row of Fig. 4 there is one wave mode with two [Fig. 4 (b)] or one [Fig. 4 (c)] bifurcations. Depending on the specific parameter values the bifurcations may be super- or subcritical, the branches may be connected or independent – this is determined by the fully nonlinear behaviour. Directly at each bifurcation there exist harmonic waves of infinitesimal amplitude with propagation speed $V_{c}=\text{Im}(\lambda)/k_{c}$ and wavelength $L_{c}=2\pi/k_{c}$, where $k_{c}$ corresponds to the neutral mode, i.e. $\text{Re}(\lambda)(k_{c})=0$. To gain knowledge about the branches of fully nonlinear travelling waves one has to resort to numerical methods.
Finite-amplitude travelling waves correspond to steady states in the co-moving frame that moves with the wave velocity $V$. Setting c=c(xVt)c=c(x-Vt), $\rho_{1}=\rho_{1}(x-Vt)$ and $\rho_{2}=\rho_{2}(x-Vt)$ we rewrite Eqs. (4) as a boundary value problem with periodic boundary conditions and an integral condition that controls the mean density $L^{-1}\int_{0}^{L}c(\chi)\,d\chi=c_{0}$, where $\chi=x-Vt$ is the phase of the wave in the co-moving frame. We then employ numerical path-continuation techniques [44, 45] bundled in the software package auto07p [46] to determine families of travelling waves. As starting states we use the analytically known small-amplitude harmonic waves, i.e. Eq. (4.1) with eigenvalue, eigenvector and wavenumber of a neutral mode.
@@
First, we consider the dry system ($\text{Ma}=0$) at fixed mean density $c_{0}=0.09$ (where the polar-nematic state does not yet exist) and examine the wave bifurcation where a branch of travelling waves emerges from the uniform nematic state. As determined in section 4.1, for $c_{0}>0.086$ this uniform state is unstable in a certain range of wave numbers that depends on $c_{0}$. When we fix $c_{0}$ and use the domain size $L$ as control parameter, i.e. as principal continuation parameter, we encounter two disconnected branches of travelling waves that emerge where the dispersion relation in Fig. 4 (b) has its two zero crossings. Fig. 5 shows in panel (a) the propagation speed $V$, in panel (b) the amplitude of the density field $A=\text{max}(c)-\text{min}(c)$, and in panel (c) the polar and mean nematic order parameter fields $\langle\rho_{1,2}\rangle=L^{-1}\int_{0}^{L}\rho_{1,2}(\chi),\,d\chi$ as a function of the domain size $L$. The wave with the larger [smaller] amplitude corresponds to a travelling density elevation [depression] both propagating in the uniform background, as shown in the insets of Fig. 5. When numerically integrating Eqs. (4) in time using a naive continuation scheme as described in section 3, at all $L$ the system converges to the state with the elevation (density hump, black line), as indicated by small star symbols in Fig. 5 (a) and (b). The travelling holes (red line) are never found in time simulations, which indicates that they are either unstable or linearly stable but with a very small basin of attraction.
@@
Second, we look at the dry system at the larger mean density $c_{0}=0.12$ where also the uniform polar-nematic state exists. The linear stability results then indicate that overall three types of small-amplitude travelling wave states may emerge from the uniform states (one from purely nematic and two from mixed polar-nematic). The corresponding zero crossings of the dispersion relations (Fig. 4) correspond to three wave bifurcations and may give rise to up to three disconnected branches of finite-amplitude waves (the exact number depends on the fully nonlinear behaviour). At $c_{0}=0.12$ we find three disconnected branches, see in Figs. 6 (a)-(c). The first one bifurcates at the critical $k$ from the purely nematic state (red dashed line) and the other two at the critical $k$’s for the mixed polar-nematic state (solid black and dot-dashed blue line). The dot-dashed branch undergoes two saddle-node bifurcations, however, at large domain sizes three clearly distinguished travelling waves remain. They correspond to a single-hump wave (solid black line) as shown in Fig. 6(1,3), a double-hump wave (blue dot-dashed line) as shown in Fig. 6(2), and a travelling-hole state (red dashed line) as shown in Fig. 6(4). The double-hump state corresponds to a large travelling density cluster followed at a small distance by a small satellite peak. Direct numerical time integration of Eqs. (4) using the naive continuation algorithm as described above only converges to single-hump waves (as indicated by star symbols in 6 (a)-(c)) indicating their linear stability. The other types of travelling waves were not found in the time simulations.
Figure 6: Three disconnected branches of finite-amplitude travelling waves for a dry system (Ma=0=0) at $c_{0}=0.12$ and $d=0.01$. (a) Wave speed $V$, (b) amplitude of the density field $A=\text{max}(c)-\text{min}(c)$ and (c) average nematic order parameter field $\langle\rho_{2}\rangle$ as a function of domain size $L$. The four red filled circles marked (1) to (4) in panel (b) indicate loci of density profiles shown in the bottom panels. Star symbols indicate stable wave states obtained by direct numerical time integration of Eqs. (4).